Two-hundred fifty people were asked whether they have seen various types of movies at a movie theater in the last month. The table gives the results of the survey. What is the experimental probability that a person has seen a comedy in the last month?
typeof movie comedy adventure animation
no who have seen type of movie 65 82 51

Answer by stanbon(75887)

65/(65+82+51) = 65/198 = 0.328...
Cheers,
Stan H.